# Bongard-RWR+: Real-World Representations of Fine-Grained Concepts in Bongard Problems

### v1

@@ Appendix C Limitations
Multiple recent studies confirm the limitations of VLMs highlighted in our work. MindSet: Vision biscione2024mindset tests models on $30$ psychological findings inspired by human visual perception, revealing fundamental differences between human and machine vision. VCog-Bench cao2024visual evaluates models on abstract and commonsense visual reasoning as well as visual question answering, showing that VLMs consistently struggle with multi-image reasoning tasks. Evaluations on the WAIS-IV test indicate that VLMs underperform in perceptual reasoning tasks galatzer2024cognitive . A systematic review of AVR performance across such benchmarks could provide a more comprehensive understanding of current model capabilities and limitations.
Appendix D Broader impacts
Appendix E Data and Code
All datasets developed in this work are publicly available via the HuggingFace Datasets platform [^3] https://huggingface.co/collections/spawlonka/bongard-rwr-plus-68252881607162f6bc884efa, released under the CC BY 4.0 License. This licensing is compatible with the terms of the Flux.1-dev model [^4] https://huggingface.co/black-forest-labs/FLUX.1-dev/blob/main/LICENSE.md, which was used to generate the images. Code for dataset generation and experimental evaluation is available on GitHub [^5] https://github.com/pavonism/bongard-rwr-plus under the MIT License.
@@ Appendix F Model prompts
Prompts $F$, and F correspond to the Describe, Augment, and Render steps in Algorithm 1, resp. Prompts F, F, F, F, and $F$ are used for the CS, I1S, D1S, I2S, and D2S tasks. Prompt F outlines the shared task introduction for both I1S and D1S, while Prompt F provides the common introduction for I2S and D2S. Prompts include $2$ illustrative examples to help models understand the task format. Importantly, these examples use different class labels to avoid biasing the model toward always selecting the same answer as in the examples. When applicable, models are instructed to respond in a structured JSON format. The validity of these outputs is enforced using the Outlines decoding backend willard2023efficient .
Prompt 1: Generate both positive and negative image descriptions. You are given an image and a general concept. Your task is to refine the concept into two concise, visually descriptive prompts: one that aligns with the image (positive prompt) and one that contrasts with it (negative prompt). Focus on making each prompt specific, clearly grounded in the image, and reflective of the core idea. You don’t need to match every detail—just convey the main visual concept. ### Example: Image: Human legs wearing socks with vertical lines Concept: Vertical lines. Positive prompt: Socks with vertical lines Negative prompt: Socks with horizontal lines Now, it’s your turn: Concept: {concept} ### Instructions: 1. Generate a positive and negative prompt based on the provided image and concept. 2. Answer using the following format: Positive prompt: Negative prompt:
Prompt 2: Augment a positive image description. You are tasked with modifying a given prompt for a diffusion model. Your primary objective is to preserve the specified **concept** while altering unrelated details or environments. You are encouraged to create diverse, creative, and unique augmentations that stay true to the concept but introduce variety in interpretation. ### Example: Prompt: "An empty white bowl with a thin black rim placed on a solid blue background." Concept: "Empty picture" Output Augmentations: [ "An empty red plate on a wooden table.", "A clear glass cup sitting on a marble countertop.", "A white ceramic vase on a patterned fabric." ] Now, it’s your turn: Prompt: {prompt} Concept: {concept} ### Instructions: 1. Generate **{n_augmentations} unique augmentations** for the provided prompt. 2. Output the results in the following JSON string array format: [ "<augmented_prompt_1>", "<augmented_prompt_2>", ... ] Ensure that each augmentation aligns with the concept and introduces creative variations in other details.
@@
Prompt 4: Concept Selection (CS) task prompt. You are a vision understanding module (an expert) designed to provide short, clear and accurate answers. The goal in solving a Bongard Problem is to identify a concept that differentiates the left and right matrix sides. You will receive a list of concepts. Only one of them correctly describes the image. Select the concept that correctly describes the image. Respond in JSON using the following format. EXAMPLE START REQUEST { "concepts": [ { "left": "Parallel lines", "right": "Perpendicular lines", "label": 1 }, { "left": "Negative slope", "right": "Positive slope", "label": 2 } ] } RESPONSE { "explanation": "The image presents several lines that don’t intersect.", "label": 1 } EXAMPLE END EXAMPLE START REQUEST { "concepts": [ { "left": "Two shapes", "right": "Six shapes", "label": 1 }, { "left": "A cycle", "right": "No cycle", "label": 2 }, { "left": "Similar shapes", "right": "Different shapes", "label": 3 }, { "left": "Small variance", "right": "Large variance", "label": 4 } ] } RESPONSE { "explanation": "All images on the left side feature circles of similar size, while the images on the right side present triangles, squares, and lines.", "label": 3 } EXAMPLE END REQUEST <request> RESPONSE
@@
Prompt 5: Shared task introduction used in both I1S and D1S settings. You are a vision understanding module (an expert) designed to provide short, clear and accurate answers. The goal in solving a Bongard Problem is to classify a test image to the corresponding class. There are two classes: LEFT and RIGHT. All images belonging to the LEFT class represent a common, shared concept which is not present in any image from the RIGHT class, and vice versa - all images belonging to the RIGHT class represent a common, shared concept which is not present in any image from the LEFT class.
@@
Prompt 6: Image-to-Side (I1S) task prompt. Solve the provided Bongard Problem. Respond in json using the following format. EXAMPLE START RESPONSE { "concept": "small vs big", "explanation": "The test image shows a small shape, similarly as all images on the left side. Conversely, the images on the right side feature big shapes.", "answer": "LEFT" } EXAMPLE END EXAMPLE START RESPONSE { "concept": "two circles vs one circle", "explanation": "The donut sits alone as the primary round element in the scene, matching the RIGHT examples where a single rounded object takes visual focus amid other details, like the soup bowl or drum.", "answer": "RIGHT" } EXAMPLE END RESPONSE
Prompt 7: Description-to-Side (D1S) task prompt. Use the descriptions below to solve the provided Bongard Problem. Respond in JSON using the following format. EXAMPLE START REQUEST { "left_descriptions": [ "A wired telephone with a coiled cord resting on a wooden table", "Power lines stretching across a clear blue sky with birds perched on them", "A clothes drying rack with colorful clothes hanging under the sunlight", "A swing hanging from a tree branch in a park with a child playing nearby", "A needle with thread passing through fabric, placed on a sewing table", "A pair of shoes with neatly tied laces placed on a tiled floor" ], "right_descriptions": [ "A smartphone with a cracked screen lying on a desk with visible damage", "A bird flying freely in the sky with its wings spread wide", "A pile of folded clothes neatly stacked on a shelf in a wardrobe", "A park bench surrounded by fallen leaves in an autumn setting", "A piece of fabric with intricate embroidery placed on a sewing table", "A pair of slip-on sandals placed on a carpeted floor near a doorway" ], "test_description": "A sailboat with its sails fully open gliding across a calm lake under a clear sky" } RESPONSE { "concept": "line is present vs no line is present", "explanation": "The first test image presents a sailboat with ropes hanging the sails. The ropes are lines, the same as items on the LEFT, like the telephone cord, power lines, and shoelaces.", "answer": "LEFT" } EXAMPLE END EXAMPLE START REQUEST { "left_descriptions": [ "Two bagels on a cutting board beside a knife", "A pair of headphones resting on a desk with the earcups facing up", "Two apples placed next to each other on a picnic blanket", "A pair of camera lenses lined up on a shelf", "A bicycle lying on the ground with both wheels visible", "A snowman with two stacked parts, one above the other" ], "right_descriptions": [ "Two forks placed on a plate", "A dartboard mounted on a brick wall with scores written on two papers beside it", "A spotlight shining onto a dark floor surrounded by cables", "A pie baking in the oven next to a tray of gingers bread men", "A record spinning on a turntable with a hand reaching toward it", "A teacup seen from above with steam rising", ], "test_description": "An octagonal dessert plate holding a single donut with sprinkles" } RESPONSE { "concept": "two circles vs one circle", "explanation": "The donut sits alone as the primary round element in the scene, matching the RIGHT examples where a single rounded object takes visual focus amid other details, like the soup bowl or drum.", "answer": "RIGHT" } EXAMPLE END REQUEST <request> RESPONSE
Prompt 8: Shared task introduction used in both I2S and D2S settings. You are a vision understanding module (an expert) designed to provide short, clear and accurate answers. The goal in solving a Bongard Problem is to classify each of the two test images to the corresponding class. There are two classes: LEFT and RIGHT. Each image belongs to exactly one class (either LEFT or RIGHT). All images belonging to the LEFT class represent a common, shared concept which is not present in any image from the RIGHT class, and vice versa - all images belonging to the RIGHT class represent a common, shared concept which is not present in any image from the LEFT class. The two test images belong to different classes.
Prompt 9: Images-to-Sides (I2S) task prompt. Solve the provided Bongard Problem. Respond in json using the following format. EXAMPLE START RESPONSE { "concept": "small vs big", "first": { "explanation": "The test image shows a small shape, similarly as all images on the left side. Conversely, the images on the right side feature big shapes.", "answer": "LEFT" }, "second": { "explanation": "The test image shows a big shape, similarly as all images on the right. The images on the left, on the other hand, feature small shapes.", "answer": "RIGHT" } } EXAMPLE END EXAMPLE START RESPONSE { "concept": "two circles vs one circle", "first": { "explanation": "The donut sits alone as the primary round element in the scene, matching the RIGHT examples where a single rounded object takes visual focus amid other details, like the soup bowl or drum.", "answer": "RIGHT" }, "second": { "explanation": "The tennis net and two balls represent two circular elements in the scene, aligning with the LEFT examples where pairs of circular objects are present, such as the bagels, apples, or camera lenses.", "answer": "LEFT" } } EXAMPLE END RESPONSE
@@
Prompt 10: Descriptions-to-Sides (D2S) task prompt. (1/2) Use the image and the descriptions below to solve the provided Bongard Problem. Respond in JSON using the following format. EXAMPLE START REQUEST { "left_descriptions": [ "A wired telephone with a coiled cord resting on a wooden table", "Power lines stretching across a clear blue sky with birds perched on them", "A clothes drying rack with colorful clothes hanging under the sunlight", "A swing hanging from a tree branch in a park with a child playing nearby", "A needle with thread passing through fabric, placed on a sewing table", "A pair of shoes with neatly tied laces placed on a tiled floor" ], "right_descriptions": [ "A smartphone with a cracked screen lying on a desk with visible damage", "A bird flying freely in the sky with its wings spread wide", "A pile of folded clothes neatly stacked on a shelf in a wardrobe", "A park bench surrounded by fallen leaves in an autumn setting", "A piece of fabric with intricate embroidery placed on a sewing table", "A pair of slip-on sandals placed on a carpeted floor near a doorway" ], "first": "A sailboat with its sails fully open gliding across a calm lake under a clear sky", "second": "A launch boat with an engine cruising through the water leaving a trail of waves behind" } RESPONSE { "concept": "line is present vs no line is present", "first": { "explanation": "The first test image presents a sailboat with ropes hanging the sails. The ropes are lines, the same as items on the LEFT, like the telephone cord, power lines, and shoelaces.", "answer": "LEFT" }, "second": { "explanation": "The second image shows a motorized launch boat, which typically doesn’t have sails or ropes visible. This aligns more with the RIGHT side examples where no lines are present, like the smartphone, slip-on sandals, or park bench.", "answer": "RIGHT" } } EXAMPLE END
@@
Prompt 11: Descriptions-to-Sides (D2S) task prompt. (2/2) EXAMPLE START REQUEST { "left_descriptions": [ "Two bagels on a cutting board beside a knife", "A pair of headphones resting on a desk with the earcups facing up", "Two apples placed next to each other on a picnic blanket", "A pair of camera lenses lined up on a shelf", "A bicycle lying on the ground with both wheels visible", "A snowman with two stacked parts, one above the other" ], "right_descriptions": [ "Two forks placed on a plate", "A dartboard mounted on a brick wall with scores written on two papers beside it", "A spotlight shining onto a dark floor surrounded by cables", "A pie baking in the oven next to a tray of gingers bread men", "A record spinning on a turntable with a hand reaching toward it", "A teacup seen from above with steam rising", ], "first": "An octagonal dessert plate holding a single donut with sprinkles", "second": "A tennis net with two balls lying nearby" } RESPONSE { "concept": "two circles vs one circle", "first": { "explanation": "The donut sits alone as the primary round element in the scene, matching the RIGHT examples where a single rounded object takes visual focus amid other details, like the soup bowl or drum.", "answer": "RIGHT" }, "second": { "explanation": "The tennis net and two balls represent two circular elements in the scene, aligning with the LEFT examples where pairs of circular objects are present, such as the bagels, apples, or camera lenses.", "answer": "LEFT" } } EXAMPLE END REQUEST <request> RESPONSE

### v2

@@ Appendix D Limitations and future work
Multiple recent studies confirm the limitations of VLMs highlighted in our work. MindSet: Vision (Biscione et al., 2024) tests models on $30$ psychological findings inspired by human visual perception, revealing fundamental differences between human and machine vision. VCog-Bench (Cao et al., 2025) evaluates models on abstract and commonsense visual reasoning as well as visual question answering, showing that VLMs consistently struggle with multi-image reasoning tasks. Evaluations on the WAIS-IV test indicate that VLMs underperform in perceptual reasoning tasks (Galatzer-Levy et al., 2024). A systematic review of AVR performance across such benchmarks could provide a more comprehensive understanding of current model capabilities and limitations.
Our evaluation of bias and fairness (Appendix E) reveals that, although Flux.1-dev offers certain demographic variation, it disproportionately generates White figures and tends to produce racially homogeneous crowds. These results are consistent with prior studies on demographic biases in Text-to-Image models (Shukla et al., 2025), highlighting the need for further research on diversity and fairness in Text-to-Image generation. Nevertheless, we do not expect these demographic imbalances to affect the benchmark’s validity for its primary purpose: evaluating abstract visual reasoning. The concepts underlying Bongard-RWR+ (e.g., spatial relations, counting, rotations) are independent of demographic attributes such as gender, race, or age. Human figures serve only as one possible way for expressing an abstract rule, not as entities whose demographic characteristics are relevant to the task.
Although we consider the expansion of Bongard-RWR dataset ($60$ instances) to Bongard-RWR+ ($5400$ instances) a significant step forward, some Bongard-RWR matrices and many original Bongard problems remain unrepresented in the extended Bongard-RWR+ dataset. As we note in Appendix E.2, this is mainly due to the limited number of images passing review and the difficulty of fully capturing complex synthetic concepts in the real-world images (e.g., "spiral curls clockwise/counterclockwise"). We believe that expanding Bongard-RWR+ to cover the missing concepts is a challenging yet promissing research direction, and we encourage the community to contribute to this endeavor. Below, we outline several directions that look promising.
(1) Advancing T2I models. In our current pipeline, human verification served primarily as a quality control step. In practice, we filtered approximately $30.2\%$ of generated images, which indicates that the majority of outputs were already acceptable. This suggests that the pipeline could operate with minimal human intervention or potentially in a fully automated manner as generative models become more reliable at following complex prompts and avoiding rendering unintended concepts. Our approach is therefore forward-compatible with future advances in generative modeling.
(2) Reducing demographic bias. The demographic skew in our dataset (Appendix E) stems from the underlying T2I model. While our pipeline does not amplify these biases, future refinements could mitigate them directly. One approach is to explicitly encode demographic diversity into prompts during both the Augment and Generate stages – for example, by specifying varied gender, racial, or age attributes when appropriate for expressing an abstract concept. Additionally, greater diversity could be achieved by leveraging a broader set of I2T, T2T, and T2I models within the pipeline, thereby reducing reliance on any single generative model.
(3) Cycle-consistency verification. In our current pipeline, a T2I model generates images from textual descriptions, and humans verify whether the outputs faithfully depict the intended concept. A natural extension is to introduce an automated cycle-consistency check: after generating an image from a source caption, an I2T model could re-caption the image to produce a reference caption. A T2T model would then compare the source and reference captions to assess semantic alignment. Images whose re-captioned descriptions fail to preserve key elements of the source prompt could be filtered automatically.
(4) Property-based testing. A more structured direction for automation is to move beyond free-form captions and operate on explicit image properties. During the Augment stage, the T2T model could be prompted not only to produce a natural language description of the scene, but also output a set of properties that the generated image should satisfy to faithfully express the intended concept. An initial closed vocabulary could include attributes such as size, shape, count, rotation, lighting, camera perspective, color, and distance. Alternatively, the model could generate open-vocabulary, instance-specific properties tailored to each concept. After image generation, an I2T model could verify whether these properties are present in the output. Images that fail to meet the specified property set (or fall below a chosen threshold) could be filtered automatically.
We note, however, that while such model-based approaches could improve scalability, some human oversight will likely remain important in the near term to ensure high data quality.
Appendix E Bias and Fairness
E.1 Demographic bias and diversity
(a) “A man standing with a surfboard…” was augmented to “A woman standing with a surfboard…”.
(b) “A woman wearing retro glasses with a pink outline…” was augmented to “A child wearing square glasses with a yellow outline…”.
(c) “A hearty bowl of Italian minestrone soup…” was augmented to “A bowl of Thai green curry with jasmine rice…”.
(d) “A herd of deer standing together in a dense forest…” was augmented to “A group of kangaroos hopping closely in a grassy plain…”.
Figure 19: Examples of (a) Gender, (b) Age, (c) Culture and (d) Geographic diversity augmentations used in Algorithm 1.
Our data generation pipeline includes several design choices aimed at promoting visual diversity and thereby mitigating potential biases. First, augmentations made in Algorithm 1 introduce variation in visual scenes, thus increasing conceptual and demographic diversity. For example, a description originally referencing “A man standing with a surfboard” was augmented to “A woman standing with a surfboard” (Fig. 19). Second, in Algorithm 2 we explicitly encourage visual diversity within BPs representing the same concept. This is accomplished through iteratively selecting image subsets that maximize intra-set visual diversity, as measured by the total pairwise cosine similarity of ViT-L/14 embeddings. Furthermore, after each matrix is constructed, we randomly exclude one image from the selected subset from future sampling, thus increasing the total number of unique images used.
Despite these methodological safeguards, Bongard-RWR+ may still inherit biases from the underlying models used in the generation pipeline, particularly Flux.1-dev for Text-to-Image generation. Since these models are trained on large-scale web data, they may reflect social, cultural, or geographic biases present in that data. To better understand these risks, we conducted a limited-scale bias audit. Two expert annotators independently reviewed a subset of $800$ images from Bongard-RWR+ depicting humans, and to each of them assigned labels related to Gender (Female, Male, Hard to tell), Race (Asian, Black, White, Hard to tell), Age (Adult, Child, Hard to tell), and Image type (Single, Pair, Crowd, Hands, Legs, Other). We assessed the annotators agreement computing Cohen’s $\kappa$ for each of the categories (Gender = $0.82$, Race = $0.84$ , Age = $0.78$, Image type = $0.89$).
The performed analysis (Fig. 20) revealed that Flux.1-dev produces a relatively balanced Gender distribution ($45.1\%$ Female, $36.6\%$ Male, $33.0\%$ Hard to tell). However, the model shows racial imbalance: $79.9\%$ of generated humans were annotated as White, rising to $98.7\%$ in images limited to human hands. Adults were far more frequent ($67.9\%$) than Children ($12.8\%$). Crowd scenes were largely racially homogeneous, with $82.2\%$ appearing to consist of a single race.
Figure 20: Distribution of the four selected human categories in Bongard-RWR+ images (Gender, Race, Age, Image type). The blue dashed line indicates the expected value under equal group representation. H2T is used as an abbreviation for Hard to tell.
E.2 Annotation reliability and bias
In our pipeline (Fig. 3), the human role is purely discriminative rather than generative and limited to filtering generated images. This substantially reduces the manual burden compared to annotation-heavy approaches such as object detection or image segmentation. Nevertheless, this dicriminative step may potentially introduce some bias. To mitigate this risk, each image generated by Algorithm 1 was independently reviewed by two expert annotators, who assigned one of the $3$ labels: Left, Right or None. This way, each annotator verified whether an image (1) adhered to the concept of its side of the matrix and (2) did not adhere to the concept of the opposing side. Images failing either criterion were flagged for exclusion.
Table 16: Percentage of discarded generated images per concept group.
Concept Group Rejection Rate
Rotation $39\%$
Contour $30\%$
Count $30\%$
Shape $27\%$
Position $26\%$
Size $19\%$
Angle $19\%$
Branching $14\%$
Similarity $14\%$
Inter-annotator agreement reached a Cohen’s $\kappa$ of $0.64$, indicating substantial consensus. Annotators disagreed on side assignment (Left vs. Right) in $5.0\%$ of the cases, disagreed between None vs. either Left or Right in $17.7\%$, and agreed on the same label in $77.3\%$. Disagreements were resolved through discussion. In case of the lack of an agreement, the image was discarded. Overall, $30.2\%$ generated images were discarded. On average, we removed $52\%$ of images for concepts that were ultimately rejected, compared to $23\%$ of them for concepts that were retained. Table 16 reports the percentage of generated images discarded during human filtering, broken down by concept group. Rotation concepts show the highest rejection rate ($39\%$), followed by Contour and Count (both $30\%$), indicating these are most challenging for the T2I model to render faithfully. In contrast, Branching and Similarity have the lowest rejection rates ($14\%$), suggesting they are easier for the pipeline to capture reliably. Overall, this strict filtering also explains why some concepts (e.g., “spiral curls clockwise/counterclockwise”) did not reach the target of $100$ matrices, due to the limited number of images passing review.
Additionally, we performed a more detailed per-concept analysis, focusing on the top five concepts with the lowest and highest image dropout. Concepts requiring strong object recognition skills had the lowest dropout (under $5\%$), such as: “A sharp projection vs. No sharp projection”, “Triangles vs. Other figures”, “Three parts vs. Four parts”, “Vertical hatched lines vs. Horizontal hatched lines”, and “Triangle inside of the circle vs. Circle inside of the triangle”. To understand why images were rejected, annotators explained their decisions for the highest-dropout concepts. The main reasons for rejection can be grouped into three categories: (1) the image did not satisfy the concept constraints, (2) the image was ambiguous (e.g., conflicting cues in the foreground and background), or (3) the image did not provide enough information to judge the concept. For example, in “Ends of the curve are parallel vs. Ends of the curve are perpendicular”, images were rejected mainly because at least one end of the curve lay outside the frame, preventing clear recognition of the concept. Selected examples are presented in Fig. 21.
(a) Left: Four parts. Right: Five parts.
(b) Left: Ends of the curve are parallel. Right: Ends of the curve are perpendicular.
(c) Left: One line. Right: Two lines.
(d) Left: Convex figures. Right: Nonconvex figures.
Figure 21: Examples of images dropped during review. (a) Depicts six petals, and therefore does not fit either side of the problem. (b) The ends of the chain extend beyond the image, making the concept ambiguous. (c) In front of the car there is a single white line, but in the background two yellow lines are visible. (d) The blood vessel appears circular, which implies convexity; however, when viewed in three dimensions, it is concave.
@@ Appendix F Model Prompts
Prompts $G$, and G correspond to the Describe, Augment, and Render steps in Algorithm 1, resp. Prompts G, G, $G$, G, and $G$ are used for the CS, I1S, D1S, I2S, and D2S tasks. Prompt G outlines the shared task introduction for both I1S and D1S, while Prompt G provides the common introduction for I2S and D2S. Prompts include $2$ illustrative examples to help models understand the task format. Importantly, these examples use different class labels to avoid biasing the model toward always selecting the same answer as in the examples. When applicable, models are instructed to respond in a structured JSON format. The validity of these outputs is enforced using the Outlines decoding backend (Willard and Louf, 2023).
Appendix G The Use of Large Language Models
We used LLMs solely to assist with polishing the manuscript, limited to grammar correction and stylistic refinement. No model contributed to research ideas, analysis, or conclusions. In our experimental pipeline, however, VLMs, LLMs, and T2I models served as the primary tools for data generation. Specifically, they were employed for (1) producing image descriptions, (2) augmenting captions, and (3) generating synthetic images. In addition, we evaluated the AVR abilities of several VLMs and LLMs.
Prompt 1: Generate both positive and negative image descriptions. {internallinenumbers*} You are given an image and a general concept. Your task is to refine the concept into two concise, visually descriptive prompts: one that aligns with the image (positive prompt) and one that contrasts with it (negative prompt). Focus on making each prompt specific, clearly grounded in the image, and reflective of the core idea. You don’t need to match every detail—just convey the main visual concept. ### Example: Image: Human legs wearing socks with vertical lines Concept: Vertical lines. Positive prompt: Socks with vertical lines Negative prompt: Socks with horizontal lines Now, it’s your turn: Concept: {concept} ### Instructions: {internallinenumbers*} 1. Generate a positive and negative prompt based on the provided image and concept. 2. Answer using the following format: Positive prompt: Negative prompt:
Prompt 2: Augment a positive image description. You are tasked with modifying a given prompt for a diffusion model. {internallinenumbers*} Your primary objective is to preserve the specified **concept** while altering unrelated details or environments. {internallinenumbers*} You are encouraged to create diverse, creative, and unique augmentations that stay true to the concept but introduce variety in interpretation. ### Example: {internallinenumbers*} Prompt: "An empty white bowl with a thin black rim placed on a solid blue background." Concept: "Empty picture" Output Augmentations: [ "An empty red plate on a wooden table.", "A clear glass cup sitting on a marble countertop.", "A white ceramic vase on a patterned fabric." ] Now, it’s your turn: Prompt: {prompt} Concept: {concept} ### Instructions: 1. Generate **{n_augmentations} unique augmentations** for the provided prompt. 2. Output the results in the following JSON string array format: [ "<augmented_prompt_1>", "<augmented_prompt_2>", ... ] {internallinenumbers*} Ensure that each augmentation aligns with the concept and introduces creative variations in other details.
@@ Appendix G The Use of Large Language Models
Prompt 4: Concept Selection (CS) task prompt. {internallinenumbers*} You are a vision understanding module (an expert) designed to provide short, clear and accurate answers. The goal in solving a Bongard Problem is to identify a concept that differentiates the left and right matrix sides. You will receive a list of concepts. Only one of them correctly describes the image. {internallinenumbers*} Select the concept that correctly describes the image. Respond in JSON using the following format. EXAMPLE START REQUEST { "concepts": [ { "left": "Parallel lines", "right": "Perpendicular lines", "label": 1 }, { "left": "Negative slope", "right": "Positive slope", "label": 2 } ] } RESPONSE { "explanation": "The image presents several lines that don’t intersect.", "label": 1 } EXAMPLE END EXAMPLE START REQUEST { "concepts": [ { "left": "Two shapes", "right": "Six shapes", "label": 1 }, { "left": "A cycle", "right": "No cycle", "label": 2 }, { "left": "Similar shapes", "right": "Different shapes", "label": 3 }, { "left": "Small variance", "right": "Large variance", "label": 4 } ] } RESPONSE { {internallinenumbers*} "explanation": "All images on the left side feature circles of similar size, while the images on the right side present triangles, squares, and lines.", "label": 3 } EXAMPLE END REQUEST <request> RESPONSE
@@
Prompt 5: Shared task introduction used in both I1S and D1S settings. {internallinenumbers*} You are a vision understanding module (an expert) designed to provide short, clear and accurate answers. The goal in solving a Bongard Problem is to classify a test image to the corresponding class. There are two classes: LEFT and RIGHT. All images belonging to the LEFT class represent a common, shared concept which is not present in any image from the RIGHT class, and vice versa - all images belonging to the RIGHT class represent a common, shared concept which is not present in any image from the LEFT class.
@@
Prompt 6: Image-to-Side (I1S) task prompt. Solve the provided Bongard Problem. Respond in json using the following format. EXAMPLE START RESPONSE { "concept": "small vs big", {internallinenumbers*} "explanation": "The test image shows a small shape, similarly as all images on the left side. Conversely, the images on the right side feature big shapes.", "answer": "LEFT" } EXAMPLE END EXAMPLE START RESPONSE { "concept": "two circles vs one circle", {internallinenumbers*} "explanation": "The donut sits alone as the primary round element in the scene, matching the RIGHT examples where a single rounded object takes visual focus amid other details, like the soup bowl or drum.", "answer": "RIGHT" } EXAMPLE END RESPONSE
@@
Prompt 7: Description-to-Side (D1S) task prompt. (1/2) {internallinenumbers*} Use the descriptions below to solve the provided Bongard Problem. Respond in JSON using the following format. EXAMPLE START REQUEST { "left_descriptions": [ "A wired telephone with a coiled cord resting on a wooden table", {internallinenumbers*} "Power lines stretching across a clear blue sky with birds perched on them", {internallinenumbers*} "A clothes drying rack with colorful clothes hanging under the sunlight", {internallinenumbers*} "A swing hanging from a tree branch in a park with a child playing nearby", {internallinenumbers*} "A needle with thread passing through fabric, placed on a sewing table", "A pair of shoes with neatly tied laces placed on a tiled floor" ], "right_descriptions": [ {internallinenumbers*} "A smartphone with a cracked screen lying on a desk with visible damage", "A bird flying freely in the sky with its wings spread wide", "A pile of folded clothes neatly stacked on a shelf in a wardrobe", "A park bench surrounded by fallen leaves in an autumn setting", "A piece of fabric with intricate embroidery placed on a sewing table", "A pair of slip-on sandals placed on a carpeted floor near a doorway" ], {internallinenumbers*} "test_description": "A sailboat with its sails fully open gliding across a calm lake under a clear sky" } RESPONSE { "concept": "line is present vs no line is present", {internallinenumbers*} "explanation": "The first test image presents a sailboat with ropes hanging the sails. The ropes are lines, the same as items on the LEFT, like the telephone cord, power lines, and shoelaces.", "answer": "LEFT" } EXAMPLE END
@@
Prompt 8: Description-to-Side (D1S) task prompt. (2/2) EXAMPLE START REQUEST { "left_descriptions": [ "Two bagels on a cutting board beside a knife", "A pair of headphones resting on a desk with the earcups facing up", "Two apples placed next to each other on a picnic blanket", "A pair of camera lenses lined up on a shelf", "A bicycle lying on the ground with both wheels visible", "A snowman with two stacked parts, one above the other" ], "right_descriptions": [ "Two forks placed on a plate", {internallinenumbers*} "A dartboard mounted on a brick wall with scores written on two papers beside it", "A spotlight shining onto a dark floor surrounded by cables", "A pie baking in the oven next to a tray of gingers bread men", "A record spinning on a turntable with a hand reaching toward it", "A teacup seen from above with steam rising", ], {internallinenumbers*} "test_description": "An octagonal dessert plate holding a single donut with sprinkles" } RESPONSE { "concept": "two circles vs one circle", {internallinenumbers*} "explanation": "The donut sits alone as the primary round element in the scene, matching the RIGHT examples where a single rounded object takes visual focus amid other details, like the soup bowl or drum.", "answer": "RIGHT" } EXAMPLE END REQUEST <request> RESPONSE
Prompt 9: Shared task introduction used in both I2S and D2S settings. {internallinenumbers*} You are a vision understanding module (an expert) designed to provide short, clear and accurate answers. The goal in solving a Bongard Problem is to classify each of the two test images to the corresponding class. There are two classes: LEFT and RIGHT. Each image belongs to exactly one class (either LEFT or RIGHT). All images belonging to the LEFT class represent a common, shared concept which is not present in any image from the RIGHT class, and vice versa - all images belonging to the RIGHT class represent a common, shared concept which is not present in any image from the LEFT class. The two test images belong to different classes.
Prompt 10: Images-to-Sides (I2S) task prompt. Solve the provided Bongard Problem. Respond in json using the following format. EXAMPLE START RESPONSE { "concept": "small vs big", "first": { {internallinenumbers*} "explanation": "The test image shows a small shape, similarly as all images on the left side. Conversely, the images on the right side feature big shapes.", "answer": "LEFT" }, "second": { {internallinenumbers*} "explanation": "The test image shows a big shape, similarly as all images on the right. The images on the left, on the other hand, feature small shapes.", "answer": "RIGHT" } } EXAMPLE END EXAMPLE START RESPONSE { "concept": "two circles vs one circle", "first": { {internallinenumbers*} "explanation": "The donut sits alone as the primary round element in the scene, matching the RIGHT examples where a single rounded object takes visual focus amid other details, like the soup bowl or drum.", "answer": "RIGHT" }, "second": { {internallinenumbers*} "explanation": "The tennis net and two balls represent two circular elements in the scene, aligning with the LEFT examples where pairs of circular objects are present, such as the bagels, apples, or camera lenses.", "answer": "LEFT" } } EXAMPLE END RESPONSE
Prompt 11: Descriptions-to-Sides (D2S) task prompt. (1/2) {internallinenumbers*} Use the image and the descriptions below to solve the provided Bongard Problem. Respond in JSON using the following format. EXAMPLE START REQUEST { "left_descriptions": [ "A wired telephone with a coiled cord resting on a wooden table", {internallinenumbers*} "Power lines stretching across a clear blue sky with birds perched on them", {internallinenumbers*} "A clothes drying rack with colorful clothes hanging under the sunlight", {internallinenumbers*} "A swing hanging from a tree branch in a park with a child playing nearby", {internallinenumbers*} "A needle with thread passing through fabric, placed on a sewing table", "A pair of shoes with neatly tied laces placed on a tiled floor" ], "right_descriptions": [ {internallinenumbers*} "A smartphone with a cracked screen lying on a desk with visible damage", "A bird flying freely in the sky with its wings spread wide", "A pile of folded clothes neatly stacked on a shelf in a wardrobe", "A park bench surrounded by fallen leaves in an autumn setting", "A piece of fabric with intricate embroidery placed on a sewing table", "A pair of slip-on sandals placed on a carpeted floor near a doorway" ], {internallinenumbers*} "first": "A sailboat with its sails fully open gliding across a calm lake under a clear sky", {internallinenumbers*} "second": "A launch boat with an engine cruising through the water leaving a trail of waves behind" } RESPONSE { "concept": "line is present vs no line is present", "first": { {internallinenumbers*} "explanation": "The first test image presents a sailboat with ropes hanging the sails. The ropes are lines, the same as items on the LEFT, like the telephone cord, power lines, and shoelaces.", "answer": "LEFT" }, "second": { {internallinenumbers*} "explanation": "The second image shows a motorized launch boat, which typically doesn’t have sails or ropes visible. This aligns more with the RIGHT side examples where no lines are present, like the smartphone, slip-on sandals, or park bench.", "answer": "RIGHT" } } EXAMPLE END
Prompt 12: Descriptions-to-Sides (D2S) task prompt. (2/2) EXAMPLE START REQUEST { "left_descriptions": [ "Two bagels on a cutting board beside a knife", "A pair of headphones resting on a desk with the earcups facing up", "Two apples placed next to each other on a picnic blanket", "A pair of camera lenses lined up on a shelf", "A bicycle lying on the ground with both wheels visible", "A snowman with two stacked parts, one above the other" ], "right_descriptions": [ "Two forks placed on a plate", {internallinenumbers*} "A dartboard mounted on a brick wall with scores written on two papers beside it", "A spotlight shining onto a dark floor surrounded by cables", "A pie baking in the oven next to a tray of gingers bread men", "A record spinning on a turntable with a hand reaching toward it", "A teacup seen from above with steam rising", ], {internallinenumbers*} "first": "An octagonal dessert plate holding a single donut with sprinkles", "second": "A tennis net with two balls lying nearby" } RESPONSE { "concept": "two circles vs one circle", "first": { {internallinenumbers*} "explanation": "The donut sits alone as the primary round element in the scene, matching the RIGHT examples where a single rounded object takes visual focus amid other details, like the soup bowl or drum.", "answer": "RIGHT" }, "second": { {internallinenumbers*} "explanation": "The tennis net and two balls represent two circular elements in the scene, aligning with the LEFT examples where pairs of circular objects are present, such as the bagels, apples, or camera lenses.", "answer": "LEFT" } } EXAMPLE END REQUEST <request> RESPONSE